A jar has twice as many red marbles as white marbles and the ratio of blue marbles to white marbles in the jar is 3:5. There are only red, white and blue marbles in the jar. If there are 2100 more red marbles than blue ones, how many marbles are in the jar altogether?
so far I came up with 6600 white, 11000 blue and 13200 red but that gives me 2200 more red and I need 2100. I just cant think anymore.. Am I close? Maria

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

"twice as many red marbles as white marbles" __ r=2w

"ratio of blue marbles to white marbles in the jar is 3:5" __ b/w=3/5 __ 5b/3=w

"2100 more red marbles than blue ones" __ r=b+2100

substituting __ b+2100=2(5b/3) __ 3b+6300=10b __ 6300=7b __ 900=b

substituting __ r=900+2100 __ r=3000

substituting __ 3000=2w __ 1500=w

Answer by stanbon(75887)

A jar has twice as many red marbles as white marbles
Eq: r -2w + 0 = 0
----------------------
and the ratio of blue marbles to white marbles in the jar is 3:5
Eq: 0 - 5w + 3b = 0
----------------------
There are only red, white and blue marbles in the jar.
If there are 2100 more red marbles than blue ones
Eq: r + 0 -b = 2100
---------------------
how many marbles are in the jar altogether?
-------------------
You have three equations with three unknows:
Solving the set of equations you would get:
r = 12600 red marbles
w = 6300 white marbles
b = 10500 blue marbles
==============
Cheers,
Stan H.